Find the Laplace Transform of half-wave and full-wave

rectified sine wave given in the following figures.

Take w = 2


1
Expert's answer
2021-05-24T16:08:42-0400

"f(t)=sin\omega t" ,


For full wave-


"L[sin\omega t]=\dfrac{1}{1-e^{-\frac{2\pi i}{\omega}}} \int_0^{2\pi} e^{-st} sin\omega t dt"


        "=\dfrac{1}{1-e^{-\frac{2\pi i}{\omega}}} (\dfrac{e^{-st}}{s^2+\omega^2}[-s sin\omega t -\omega cos\omega t])_0^{\frac{\pi}{\omega}}"

 

        "=\dfrac{1}{1-e^{-\frac{2\pi i}{\omega}}} (\dfrac{e^{-\frac{s\pi}{\omega}} \omega+\omega}{s^2+\omega^2})"


       "=\dfrac{\omega}{(1-e^{-\frac{\pi i}{\omega}} )(s^2+\omega^2)}"


For half wave-


"L[sin\omega t]=\dfrac{1}{1-e^{-\frac{2\pi i}{\omega}}} \int_0^{\pi} e^{-st} sin\omega t dt"


        "=\dfrac{1}{1-e^{-\frac{2\pi i}{\omega}}} (\dfrac{e^{-st}}{s^2+\omega^2}[-s sin\omega t -\omega cos\omega t])_0^{\frac{\pi}{2\omega}}"

 

        "=\dfrac{1}{1-e^{-\frac{\pi i}{\omega}}} (\dfrac{e^{-\frac{s\pi}{2\omega}} \omega+\omega}{s^2+\omega^2})"


       "=\dfrac{2\omega}{(1-e^{-\frac{\pi i}{\omega}} )(s^2+\omega^2)}"